OpenStudy (anonymous):

Which line shows the first error in the solution? Four less than a number divided by 3 is __11. What is the number? @SolomonZelman

OpenStudy (solomonzelman):

are you sure there is an eorr?

OpenStudy (anonymous):

in the link i copied

OpenStudy (solomonzelman):

does "Four less than a number divided by 3 is __11" is where the work in your link is coming from, or am I looking at the 1st line, as at the given?

OpenStudy (anonymous):

So the question is like this:

OpenStudy (anonymous):

Which line shows the first error in the solution? Four less than a number divided by 3 is __11. What is the number? ///////////////////////// ///////////////////////// ///////////////////////// ///////////////////////// A. Line 1 B. Line 2 C. Line 3 D. Line 4

OpenStudy (anonymous):

///// indicates the picture in the link

OpenStudy (solomonzelman):

and what is the ___ next to 11?

OpenStudy (solomonzelman):

is __ suposed to be the =?

OpenStudy (anonymous):

no the -

OpenStudy (anonymous):

-11

OpenStudy (solomonzelman):

minus?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Four less than a number divided by 3 is - 11

OpenStudy (solomonzelman):

so then you tell me, what the error is (in which line) if he was supposed to write: Four less than a number divided by 3 is - 11 and wrote: Four less than a number divided by 3 is = 11

OpenStudy (texaschic101):

its the problem.. should be : x/3 - 4 = -11

OpenStudy (solomonzelman):

it is "Four less than a number divided by 3 is - 11" and we are figuring which line has an error (based on the screenshot)

OpenStudy (anonymous):

yes @SolomonZelman

OpenStudy (solomonzelman):

so zkmath, can you answer my QW?

OpenStudy (solomonzelman):

(I was explaining what the problem is to others, and the question is for you)

OpenStudy (anonymous):

how do you mean

OpenStudy (solomonzelman):

I am reporting the question (below) : ~~~~~~~~~~~~~~~~~~~~~~~~~~~ so then you tell me, what the error is (in which line) ? IF he was supposed to write: Four less than a number divided by 3 is - 11 AND wrote: Four less than a number divided by 3 is = 11

OpenStudy (texaschic101):

it doesn't matter...4 less then a number means subtract

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 5 is 3 more than –6. What is the number?

OpenStudy (solomonzelman):

wait, or is it: Four less than a number divided by 3 is = - 11

OpenStudy (solomonzelman):

I am sorry, the way you gave the question confused me, so is that /\ /\ the question?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

anyways. there are two errors in 1 line then

OpenStudy (solomonzelman):

but which 1?

OpenStudy (anonymous):

Which line shows the first error in the solution? Four less than a number divided by 3 is __11. What is the number? (here must the picture be) A. Line 1 B. Line 2 C. Line 3 D. Line 4

OpenStudy (anonymous):

-11

OpenStudy (solomonzelman):

you are starting like this: line 0: "Four less than a number divided by 3 is = - 11" (and then the rest of the lines, as on the screenshot)

OpenStudy (anonymous):

there is not a line 0

OpenStudy (solomonzelman):

imagine please that there is

OpenStudy (anonymous):

okay then yes just like that

OpenStudy (texaschic101):

line 1 is your answer....the problem is written incorrectly

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

texaschic101 not patient enough?

OpenStudy (texaschic101):

but I do not know what the __ in your problem means

OpenStudy (solomonzelman):

it is minus

OpenStudy (solomonzelman):

but he is claiming it is not =-, he says it is just -.

OpenStudy (texaschic101):

so its simple : x/3 + 4 = -11 ??

OpenStudy (solomonzelman):

that is what I said it to be, but he said no

OpenStudy (solomonzelman):

wait, less than 4, -4.

OpenStudy (texaschic101):

there has to be an equal sign or you cannot do the problem

OpenStudy (texaschic101):

that is what I was saying...less then means minus

OpenStudy (texaschic101):

so the problem should be : x/3 - 4 = -11

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

yes, but you just gave away the answer, without previously explaining..... I can feel why, but next time, please...

OpenStudy (texaschic101):

ok...I understnad. I will do better.....sorry

OpenStudy (solomonzelman):

it's alright:)

OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 5 is 3 more than –6. What is the number?

OpenStudy (texaschic101):

ok....how would you write a number divided by 5 ??

OpenStudy (anonymous):

45

OpenStudy (anonymous):

-45 i mean

OpenStudy (texaschic101):

let the number be represented by x x/5 = a number divided by 5 " is " means = 3 more then -6.......-6 + 3 so we have : x/5 = -6 + 3 x/5 = - 3 --- multiply both sides by 5 x = - 15

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thanks @texaschic101

OpenStudy (texaschic101):

anytime zkmath :)

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

what do you think ?

OpenStudy (anonymous):

perfect!